OpenStudy (anonymous):

Use the formula for lowering the powers to simplify the expression: (a) If cos^2 3 x - sin^2 3 x = cos (f(x)), then f(x) = ; (b) If cos^4 3 x - sin^4 3 x = cos (g(x)), then g(x) = . For part (b), use a^4-b^4=(a^2-b^2)(a^2+b^2)!

zepdrix (zepdrix):

For part A, we'll want to recall the `Double Angle Identity for Cosine`. It shows up in several different forms. But here is the one we're interested in. \[\large \cos^2\theta-\sin^2\theta=\cos(2\theta)\] Hmm how can we maybe use this to solve part A?

OpenStudy (anonymous):

I found part a but now I just cant get part B